### v2

# BOD: Blindly Optimal Data Discovery

Abstract.
Combining discovery and augmentation is important in the era of data usage when it comes to predicting the outcome of tasks. However, having to ask the user the utility function to discover the goal to achieve the optimal small rightful dataset is not an optimal solution. The existing solutions do not make good use of this combination, hence underutilizing the data. In this paper, we introduce a novel goal-oriented framework, called BOD: Blindly Optimal Data Discovery, that involves humans in the loop and comparing utility scores every time querying in the process without knowing the utility function. This establishes the promise of using BOD: Blindly Optimal Data Discovery for modern data science solutions.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/Thomas12012002/BOD.
1. Introduction
When data scientists have multiple datasets, each contains various variables, and scientists want to know what they could get from the data, determining the outcome after joining multiple relational tables. Assume all the datasets augmenting are valid candidates, the values are integers, no overlapping values with different attributes and each has no missing number value; scientists often want to find the goal from the datasets after augmenting in which they want to know what data can do with or predicting what based on the datasets, for example housing prices are often based on location (which contains near urban center or a place where the criminal status is low) and goverment policies (which has taxes incentive) or the home values ( including the age and size of the house). To predict the housing price, scientists with their expertise in the field, rank each variable in each dataset one at a time until there is no variable in any dataset remaining and augment the datasets which are location, government policies in the area, home values.
The process of using machine learning application and causal relationship tasks to predict the outcome requires a utility function that users have to know exactly beforehand. This would be abundant for users to walk them through the process of finding utility functions based on data in datasets when users do not know their exact information. We solve this problem by introducing BOD: Blindly Optimal Data Discovery by eliminating the finding utility function process and by asking a small number of questions based on variables of datasets to output a small subset of rightful tuples, which is the goal of predicting the future.
The work of algorithms: The query first asks the user to pick the most important variable in each dataset. For example, if a user want to predict housing price without knowing their utility function, the query asks user to rank each variable in each of the datasets once at a time in multiple rounds. Based on the user’s ranking, the query selects all the tuples that meet the user’s preferences based on utility score, and the process continues working with these selected tuples. The query then outputs all the desirable tuples whenever user stops ranking the variables.
Figure 1. Augmenting data from tables and filtering out items based on human feedback of their ranking for variables
2. Related Studies
Machine Learning. To make use of data to get to the point of predicting or getting to know what the data can do, the technique of using machine learning in the work (Galhotra et al., 2023) requires the user’s utility function. However, this approach is not applicable in the case when the user’s utility function is unknown. When in the case of an unknown utility function, the work of (Gong et al., 2021) only selects the nearly optimal solution without understanding what the data can do, our algorithm, the Blindly Optimal Data Discovery, combines the advantages of these two works, which are getting to know what the data can do and unknown user’s utility function.
Top-k. Top-K algorithm has many developments and ideas for a long time (Fraternali et al., 2012) (Lee et al., 2009) (Lian and Chen, 2009) (Qin et al., 2012) (Soliman et al., 2007), however, the algorithm requires user’s utility function, which is not applicable in many cases and hard to obtain itself.
Skyline. There are many studies related to Skyline algorithm including (Lee et al., 2009) (Lin et al., 2007) (Mindolin and Chomicki, 2009) (Chan et al., 2006a) (Chan et al., 2006b) (Tao et al., 2007) (Tao et al., 2009) (Papadias et al., 2005) (Goncalves and Yidal, 2005) (Xia et al., 2008). The Skyline algorithm, the disadvantage would be the output size, which we have no control over, which means there may be uninterested tuples in the output and the output may leave out the tuples that are dominated by others even the user have interest in those. To fix this problem, researchers presented (Mouratidis et al., 2021) combining the technique of top-k and Skyline together to control the size of the output
Blindly Optimal Data Discovery. Our algorithm involves human interaction with the data to rank the variables in each of the datasets once at a time, and we filter out the promising tuples considering the utility score after each ranking phase.
3. Problem Definition
Table 1. Notation and meaning
$D$ The set of all tuples
$d$ The total attributes of the final table after augmenting all the datasets
$S$ The table of all the relational tables augmented over $n_{i}$ attributes
$t$ The specific tuple that has an instance schema
$y_{min_{i}}$ The maximum values’s summation based on user’s $i$th ranking
$y_{max_{i}}$ The summation of all values based on user’s $i$th ranking
$y_{i}$ The summation of all values of the instance
Let $(S_{1},\dots,S_{n})$ denote a relation table over $n$ attributes, $S_{i}$ denotes the $i$th attribute. $D$ composes the list of tuples with a table $(S_{1},\dots,S_{n})$ such that every tuples $t\in T$ is an instance of the table.
Linear function:
  • $LINEAR=\{f|f(x)=\sum\limits_{i=1}^{d}f(x_{i})\\ \text{\{where each}\ f(x_{i})\text{ is a linear function.}\}$. }
Assume we want to predict the housing price in an area, we have datasets including location, government policies in the area, and home values. In detail, the dataset location composes a relational table that has two attributes that are Near Urban and Criminal Free, the dataset government policies composes table that has one attribute which is Tax, while the dataset home values composes a tables that has two attributes which are Size and Age of the house. After augmenting all the datasets, we have a dataset composes five attributes and six instances after augmenting the tables as shown in the table 2
Price Near Urban Criminal Free Tax Size Age
House 1 26 5 90 1500 150
house 2 35 2 20 1300 120
House 3 45 3 78 2000 95
House 4 9 2 46 1700 50
House 5 6 4 65 1800 25
house 6 47 7 30 1450 75
Table 2. Example of the tables after augmenting
Then we scale down the values in the range of $\{0,1\}$ by the highest value in each variables and we calculate $y_{min_{i}}$ after asking the user to rank the variables in the original tables as shown in table 3.
Price Near Urban Criminal Free Tax Size Age Sum
House 1 0.55 0.71 1 0.75 1 4.01
house 2 0.74 0.29 0.22 0.65 0.8 2.7
House 3 0.96 0.43 0.87 1 0.63 3.89
House 4 0.19 0.29 0.51 0.85 0.33 2.17
House 5 0.13 0.57 0.72 0.9 0.17 2.49
house 6 1 1 0.33 0.73 0.5 3.56
Table 3. Example of the query scaling down all the values in all variables after augmenting
Suppose the user rank Near urban , Tax, and Size as the first priority in the first interaction with the algorithm, we then calculate the summation of each instance and find the maximum calculation in this round, which is $y_{min_{1}}=2.83$ in this case as shown in table 4.
Price Near Urban Tax Size Near Urban + Tax + Size
House 1 0.55 1 0.75 2.3
house 2 0.74 0.22 0.65 1.61
House 3 0.96 0.87 1 2.83
House 4 0.19 0.51 0.85 1.55
House 5 0.13 0.72 0.9 1.75
house 6 1 0.33 0.73 2.06
Table 4. Example of the query when user rank Near Urban, Tax, and Size as the first priorities
Based on the information from the first interaction, we find the tuples that have $y_{min_{1}}\leq y_{1}\leq y_{max_{1}}$ that are House 3 and House 6 since $2.83\leq y_{1}\leq 3.89$. The table 5 shown the result from the first interaction. The final interaction which is the second ranking would result in only House 3.
Price Near Urban Criminal Free Tax Size Age Sum
House 3 0.96 0.43 0.87 1 0.63 3.89
house 6 1 1 0.33 0.73 0.5 3.56
Table 5. The results of the algorithm
4. BOD: Blindly Optimal Data Discovery
Since we scale down each value in a variable by a factor of max value of the variable such that every value now in the variable is in the range of $\{0,1\}$ and we scale the values in every variable after augmenting all the datasets. For each of the relational tables, We ask the user to pick the first most important variable in each of dataset, knowing that the number of variables in each table is varied, called $n_{i}$. We then find the maximum sum of values from the most first important variables from the original tables after augmenting, called $y_{min_{1}}$. Then, we find $y_{max_{1}}$ from the row with the maximum sum $y_{min_{1}}$, which is the sum of all values in that row. After that, we select the items that have their utility score, $y_{1}$ , which is the sum of all variables, in which $y_{min_{1}}\leq y_{1}\leq y_{max_{1}}$. We then use these items in a new dataset and begin the second phase, and so on, until we get to the last variable in any table or the user wants to stop ranking the variables in each dataset.
Algorithm 1 Blindly Optimal Data Discovery
0:  The tables $S_{i}$ that has $D$ tuples.
0:  return a subset $t$ of $D$.
1:  Let $S_{i}=1$
2:  while $S_{i}\leq$ length of $S$ do
3:     for $i$ in the length of $S_{i}$ do
4:        Let the user rank the attributes in each table $S_{i}$ until the user stops
5:     $S_{i}+=1$
6:  Find the maximum length of each $S_{i}$ in $S$, called $max_{n_{i}}$
7:  Concatenate DataFrames/Tables horizontally along columns
8:  $d_{i}=1$
9:  for $d_{i}\leq d$ do
10:     find max value called $S_{m}ax_{i}$
11:     Scale down all values by dividing $S_{max_{i}}$
12:     $d_{i}+=1$
13:  Update the new $S$
14:  for $i$ in range $max_{n_{i}}$ do
15:     Based one user ranking in $i$th time, calculate $y_{min_{i}}$ and $y_{max_{i}}$ accordingly
16:     If $y_{min_{i}}\leq y_{i}\leq y_{max_{i}}$
17:     Select all tuples $t$ that satisfies and let update $S$
18:  return  $t$
5. BOD’s Performance Guarantee
We show that Blindly Optimal Data Discovery can identify the optimal solution in exactly $O(d)$ queries and returns a nearly optimal solutions in a maximum of $O(d)$ queries since the number of questions depends on the amount of the total attributes after augmenting all the datasets. The performance runs far better than any previously known algorithms.
6. Experimental Evaluation
We assume the augmentation part is correct with no error from the scientist’s domain; all values are valid, not missing, and all are in positive integers. We test the algorithm using Google Collaboration running on GL65 Leopard 10SCXK, an x64-based PC, on Microsoft Windows 11 Home Single Language. We test on three synthetic datasets, each dataset has 10000 tuples and the range of 3 to 9 variables. The values in each cell are randomly generated in the range of $\{1,1000\}$.
Figure 2. Changes in number of attributes
Then, We test on three synthetic datasets in the range of 5000, 10000, and 15000 tuples, each with six variables. The values in each cell are randomly generated in the range of $\{1,1000\}$.
Figure 3. Changes in number of tuples
7. Conclusion
The advantage is that the algorithm does not depend on the utility function and asks users the maximum amount of questions based on the total attributes after augmenting all the datasets, which is $d$. The algorithm outputs a small subset of items that help predict the optimal outcomes scientists could use with data science applications. The running time that we test in the case of three synthetic datasets, each dataset has 10000 tuples and a range of 3 to 9 variables. The values in each cell are randomly generated in the range of $\{1,1000\}$ shown in Figure 2, have the maximum running time under 30 seconds while the same maximum running time happens in the case of three synthetic datasets in the range of 5000, 10000, and 15000 tuples, each with six variables. The values in each cell are randomly generated in the range of $\{1,1000\}$.
8. Citations